A brass cube of side 10cm is heated through 30C.If the linear expansivity of brass is 2.0*10^-5K^-1 what is the increase in its volume


1
Expert's answer
2022-02-21T12:06:20-0500

Given:

"a_0=10\;\rm cm"

"\Delta T=30\:\rm K"

"\alpha=2.0*10^{-5}\: \rm K^{-1}"


The thermal volume expansion is described by an equation


"\Delta V=V_0(3\alpha)\Delta T\\""=\rm (10\: cm)^3*(3*2.0*10^{-5}\:K^{-1})*30\: K=1.8\: cm^3"